OpenStudy (anonymous):

Bernie counted the cars and motorcycles that passed his house in 45 minutes. He counted 26 vehicles and 88 tires. How many motorcycles passed his house? 7 8 9 10

OpenStudy (anonymous):

c = number of cars....m = number of motorcycles c + m = 26 Number of tires on cars + Number of tires on motorcycles = 88 4c + 2m = 88

OpenStudy (the_fizicx99):

@Hannnnah there's 2 ways to doing this, I'd recommend using what Bangkok wrote above. Do you know how to solve systems of equations?

OpenStudy (anonymous):

No this isn't my school work its my little brothers and i have to check it for him but I dont know if the answers were right so yeah I just asked on here

OpenStudy (the_fizicx99):

Mkay, but I mean do you know how to solve systems of equations?

OpenStudy (anonymous):

No I haven't done this in like 2 years lol

OpenStudy (the_fizicx99):

k, I'd use substitution method, C + M = 26 4C + 2M = 88 M = -C + 26 4C + 2(-C + 26) 4C -2C + 52 = 88 2C + 52 = 88 -52 -52 2C = 36 /2 /2 C = 18 with common logic you know that 18 cars passed by so if a total of 26 pass and 18 were cars, 26 - 18 = 8, so 8 were motorcycles. Or, C = -M + 26 4(-M + 26) + 2M = 88 -4M + 2M + 104 = 88 -2M + 104 = 88 -104 -104 -2M = -16 /-2 /-2 M = 8 M being motorcycles

OpenStudy (anonymous):

Okay thanks so much (:

OpenStudy (the_fizicx99):

Yw, "2 years ago" what grade are you in now o.o

OpenStudy (anonymous):

Im not in high school lol

OpenStudy (the_fizicx99):

Ah, was curious if they didn't teach you this before o.0

OpenStudy (anonymous):

I dont remember any of it

OpenStudy (the_fizicx99):

Oh i thought you weren't in highschool \(\ yet \)

OpenStudy (anonymous):

Nope lol

OpenStudy (the_fizicx99):

I'm lost now, I'm just going to assume you're not in high school yet :> correct me if i'm wrong

OpenStudy (anonymous):

No I am not i am not in school lol